OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)?

OpenStudy (anonymous):

This one is a little different. First you have to find your slope using the slope formula. (8-5) / (-1 - 2) Slope = 3/-3, or just -1. Now plug this into y = mx + b as m. y = -1x + b. Plug in either set of points you are given for y and x, solve for b, plug b back into y = -1x + b.

OpenStudy (anonymous):

b=7 Meanin y=-1x+7 Whats Next

OpenStudy (anonymous):

y-y1/x-x1 slope y = m(x - a) + b equation Learn.